### v1

# Are Image Distributions Indistinguishable to Humans Indistinguishable to Classifiers?

Abstract
The ultimate goal of generative models is to characterize the data distribution perfectly. For image generation, common metrics of visual quality (e.g., FID), and the truthlikeness of generated images to the human eyes seem to suggest that we are close to achieving it. However, through distribution classification tasks, we find that, in the eyes of classifiers parameterized by neural networks, the strongest diffusion models are still far from this goal. Specifically, classifiers consistently and effortlessly distinguish between real and generated images in various settings. Further, we observe an intriguing discrepancy: classifiers can identify differences between diffusion models with similar performance (e.g., U-ViT-H vs. DiT-XL), but struggle to differentiate between the smallest and largest models in the same family (e.g., EDM2-XS vs. EDM2-XXL), whereas humans exhibit the opposite tendency. As an explanation, our comprehensive empirical study suggests that, unlike humans, classifiers tend to classify images through edge and high-frequency components. We believe that our methodology can serve as a probe to understand how generative models work and inspire further thought on how existing models can be improved and how the abuse of such models can be prevented.
1 Introduction
Diffusion probabilistic models [1, 2, 3] have emerged as leading generative models for image [4, 5, 6, 7], video [8, 9, 10, 11] and 3D content generation [12, 13, 14, 15], surpassing previous generative models such as generative adversarial networks [16], variational autoencoders [17], and normalizing flows [18]. In particular, focusing on image generation, diffusion models [19, 20, 21, 22, 23, 24] can generate realistic images that are indistinguishable from real ones to the human eyes (see Fig. 2). Moreover, common quantitative metrics like FID [25] indicate that the generated distribution from state-of-the-art diffusion models is already close to the ImageNet validation set [19, 21, 20, 26, 27].
In this context, we are curious about a fundamental question: How far is the generated distribution from the real distribution? Binary classification is a natural approach, as classifiers parameterized by neural networks can serve as effective indicators of the differences between the real and generated distributions through classification accuracy. Inspired by the methodology in the work [28, 29] on dataset bias [30], we propose distribution classification tasks, predicting the distribution from which an image is taken, naturally serving as a probe to explore this critical question.
We first check are image distributions indistinguishable to humans indistinguishable to classifiers? In particular, we train commonly used classifiers to distinguish between real and generated images across various settings. To our surprise, all classifiers effortlessly achieve extremely high distribution classification accuracy with far fewer training samples than required for the diffusion models. Quantitatively, most classifiers can achieve over 98% accuracy. Furthermore, even self-supervised classifiers demonstrate the ability to identify generated images. Therefore, in the eyes of classifiers parameterized by neural networks, the generated distribution remains significantly different from the real distribution, indicating that there is still substantial room for improvement in diffusion models.
More surprisingly, we observe intriguing contradictions between classifiers and both FID and human judgments through the following two groups of experiments: The first group aims to distinguish between different diffusion models with similar FID scores and visual quality but different inductive biases. The second group aims to classify the smallest and largest models within the same diffusion model family, which share inductive biases but differ in visual quality. Notably, classifiers perform well on the first task but struggle with the second, while humans exhibit a different trend. This demonstrates that, unlike humans, classifiers focus more on inductive biases than visual quality.
Figure 1: Four-way distribution classification tasks: Which one is real in each row? [^2] From left to right, the first two rows show images from U-ViT, ImageNet (real), DiT, and EDM2-XXL. The bottom row shows images from Pixart-$\alpha$, Playground-v2.5, COCO (real), and SDXL.The samples are from real images or generated from state-of-the-art diffusion models. Notably, classifiers consistently and effortlessly distinguish between real and generated images in all settings.
Furthermore, we delve into how classifiers distinguish between different distributions to better understand their behaviours. Using Grad-CAM [31], we observe that classifiers primarily rely on edge area to distinguish real and generated images. When we center-crop images at various resolutions to remove edges, classifiers still successfully differentiate based on the new edges, even at the smallest resolution of 16. Additionally, random cropping at resolutions of 128 and 64, each performed four times, maintains high classification accuracy, indicating pervasive biases recognized by classifiers. Additionally, we observe that classifiers achieve high accuracy except when only minimal low-frequency components are present. In contrast, for models within the same diffusion family, given any specific frequency bands, classifiers’ accuracy drops close to random guessing.
Our findings have several implications. Firstly, our methodology can serve as a probe to understand how generative models work. For instance, our results indicate that current diffusion models have significant room for improvement in edge areas and high-frequency components. Enhancing these aspects will better match the data distribution. Secondly, our findings suggest that the visual quality of existing diffusion models is already sufficient for content generation, as high-frequency components are generally imperceptible to humans. Lastly, our work highlights the potential for designing more efficient algorithms based on edge areas or high-frequency components to detect generated images.
Briefly, our key contributions are as follows:
  • We find that classifiers can effortlessly distinguish diffusion-generated and real distributions, even if diffusion models achieve remarkably low FID and generate lifelike images (see Fig. 2).
  • We demonstrate that classifiers, unlike humans, are more sensitive to the inductive biases of different diffusion models but less sensitive to visual quality.
  • We explore how classifiers distinguish between real and generated images, finding that they primarily rely on edge information. They maintain high accuracy even with limited frequency components, except when only minimal low-frequency components are available.
2 Background
2.1 Diffusion model
Diffusion models [1, 2, 3] gradually inject noise into data $\bm{x}$ during the forward process:
$\bm{z}_{t}=\sqrt{\bar{\alpha}_{t}}\bm{x}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon},$ (1)
and remove noise to generate data in the reverse process. Diffusion models typically use a noise prediction network $\bm{\epsilon}_{\bm{\theta}}(\bm{z}_{t},t)$ to predict the noise $\bm{\epsilon}$ added to $\bm{z}_{t}$. Noise prediction loss is defined as:
$\mathcal{L}=\mathbb{E}_{t,\bm{x}_{0},\bm{\epsilon}}[\|\bm{\epsilon}_{\bm{\theta}}(\bm{z}_{t},t)-\bm{\epsilon}\|_{2}^{2}].$ (2)
2.2 Frequency filter
The frequency content of an image reflects the rate of pixel value changes. Low-frequency components represent overall outline and slow grayscale changes, while high-frequency components capture fine details like object edges and textures. Notably, convolutional neural networks can capture high-frequency components that are usually imperceptible to humans [32].
Low-pass filters are implemented as follows: Compute the Fourier transform of the image $f(x,y)$ using $F(u,v)=\text{FFT}(f(x,y))$, and then center the spectrum with $F_{c}(u,v)=\text{fftshift}(F(u,v))$. Define a rectangular mask $H(u,v)$ in the frequency domain, where $M$ and $N$ are the image dimensions:
$H(u,v)=\begin{cases}1&\text{if }\left|u-\frac{M}{2}\right|\leq\text{threshold}\text{ and }\left|v-\frac{N}{2}\right|\leq\text{threshold},\\ 0&\text{otherwise}.\end{cases}$ (3)
Apply the mask by multiplying $G(u,v)=F_{c}(u,v)\odot H(u,v)$, reverse the shift with $G_{c}(u,v)=\text{ifftshift}(G(u,v))$, and transform back to the spatial domain using $g(x,y)=\text{IFFT}(G_{c}(u,v))$.
High-pass filters are implemented similarly, but with the mask defined to pass high frequencies:
$H(u,v)=\begin{cases}0&\text{if }\left|u-\frac{M}{2}\right|\leq\text{threshold}\text{ and }\left|v-\frac{N}{2}\right|\leq\text{threshold},\\ 1&\text{otherwise}.\end{cases}$ (4)
Band-pass filters combine low-pass and high-pass filters, allowing frequencies between low and high thresholds to pass while setting others to 0.
3 Related work
Dataset bias. Torralba and Efros [30] introduced the "Name That Dataset" game to discover dataset bias. Tommasi et al. [29] and Liu and He [28] expanded on this by applying convolutional neural networks and large-scale datasets. Inspired by the methodologies in these works, we propose distribution classification tasks, predicting the distribution from which an image is taken. In this paper, the distribution refers to either the real distribution or the diffusion-generated distribution.
Generated image detection. We mention some empirical work on generated image detection in three key aspects: universal detection [33, 34, 35, 36], reconstruction-based methods [37, 38], and frequency analysis [32, 39, 40, 41, 42, 43, 44, 45, 46, 47]. Frequency analysis, closely related to our work, shows that generative data usually differ from real data in the frequency spectrum, particularly in high-frequency components. Although diffusion models exhibit fewer frequency artifacts compared to GANs, they still mismatch real images in high-frequency components. This finding aligns with our results. Unlike these works, which rely on frequency spectrum analysis, our approach uses distribution classification tasks to explore the differences.
4 Experimental settings
We present the main experiment settings as follows. For more details, please see Appendix A.
Dataset. For label-to-image tasks, we use the CIFAR-10 [48] and ImageNet [49] datasets, and for text-to-image tasks, we use the COCO2014 dataset [50]. For CIFAR-10, we construct the real distribution from its training and validation sets, generating 50k training images and 10k validation images using a diffusion model for the generated distribution. For ImageNet, we consider resolutions of 256 and 512. We randomly sample 100k training images and 50k validation images from ImageNet’s respective sets for the real distribution, and generate an equivalent number of images using diffusion models for the generated distribution. We follow ADM’s [51] method to preprocess ImageNet into these resolutions. For COCO2014, we randomly sample 10k training images and 1k validation images from its respective sets for the real distribution, and use the corresponding prompts to generate an equivalent number of images for the generated distribution.
Classifier. By default, we use ResNet-50 [52] as the classifier architecture, and also consider ConvNeXt-T [53] and ViT-S [54] for completeness. Our pre-processing follows standard supervised training [53]. Specifically, during training, the classifiers process randomly augmented crops of $224\times 224$ images. For validation, images are resized to 256 pixels on the shorter side, preserving the aspect ratio, and then center cropped to $224\times 224$. For CIFAR-10 experiments, we initialize the classifier with ResNet-50 pre-trained on ImageNet and also provide the results trained from scratch (see Appendix C.1). For ImageNet, we train ResNet-50 from scratch.
Diffusion model. For the generation of CIFAR-10, we consider two diffusion models: EDM [26] and U-ViT [20]. For the generation of ImageNet-256, we explore three diffusion models: EDM2 [19], U-ViT-H/2 [20], DiT-XL/2 [21]. We resize EDM2’s generated images from 512 to 256 resolution, resulting in EDM2-XXL achieving an FID of 2.14, similar to U-ViT-H/2 and DiT-XL/2. For the generation of ImageNet-512, we use EDM2-XXL [19]. For the generation of COCO, we consider three text-to-image diffusion models: Pixart-$\alpha$ [23], SDXL [24], and Playground-v2.5 [22].
Evaluation. We use the top-1 accuracy on the validation set to evaluate classification performance.
Abbreviations. In the figures and tables following this paper, the following abbreviations are used: CIFAR-10 (C), ImageNet (I), U-ViT (U), EDM (E), DiT-XL/2 (D).
5 Generated images are easily classified as fake
State-of-the-art diffusion models achieve remarkably low FID scores and can generate lifelike images (see Fig. 2). These results suggest that diffusion models are approaching the ability to perfectly learn the distribution of real images. Given this context, our initial series of experiments aims to determine whether classifiers can distinguish diffusion-generated and real distribution.
5.1 Classifiers achieve high accuracy across various settings
Label-to-image. As shown in Tab. 1, we present binary distribution classification accuracy on various combinations of real and generated distributions in the label-to-image scenario. To our surprise, in the eyes of classifiers defined by the neural network, the generated distribution is still significantly different from the real distribution. Specifically, all classifiers are capable of effortlessly distinguishing between real and generated images across various dataset combinations. Quantitatively, most classifiers can achieve over $>98\%$ distribution classification accuracy.
Moreover, we find a positive correlation between distribution classification accuracy and the number of training samples per distribution of the combination of U-ViT-H/2 and ImageNet-256 (see Fig. 2). Notably, using only 50k samples per distribution (less than 10% of the number of ImageNet samples, which the generator is trained on), classifiers achieve an accuracy exceeding 99.5%. Besides, the clear positive correlation suggests that classifiers are learning more generalizable features as the sample size increases rather than simply memorizing samples.
Table 1: Binary distribution classification on label-to-image. All classifiers yield high accuracy on various datasets against strong generative models. FIDs are taken from the corresponding references.
Real dataset Generative model Corrresponding FID Classifier Accuracy (%)
C U [20] 3.11 ResNet-50 99.92
ViT-S 98.04
ConvNeXt-T 99.96
C E [26] 1.79 ResNet-50 96.25
ViT-S 89.38
ConvNeXt-T 98.43
I-256 U-H/2 [20] 2.29 ResNet-50 99.95
ViT-S 98.06
ConvNeXt-T 99.89
I-512 E2-XXL [19] 1.81 ResNet-50 99.73
ViT-S 95.15
ConvNeXt-T 86.44
Table 2: Four-way distribution classification on text-to-image. All classifiers yield high accuracy in distinguishing four distributions: COCO [50], Pixart-$\alpha$ [23], SDXL [24], and Playground-v2.5 [22], using only 5k or 10k training samples per dataset. Notably, as the number of training samples increases, the accuracy of distribution classification consistently improves.
Training samples Model Accuracy (%)
5k ResNet-50 88.28
10k ResNet-50 93.45
5k ViT-S 76.33
10k ViT-S 83.53
5k ConvNeXt-T 80.53
10k ConvNeXt-T 85.13
Figure 2: Positive correlation between accuracy and number of training samples per distribution. With only 50k samples per distribution, classifiers achieve over 99.5% accuracy.
Text-to-image. As shown in Tab. 2, we present the distribution classification accuracy in the text-to-image scenario. We consider combinations of four distributions: COCO [50], Pixart-$\alpha$ [23], SDXL [24], and Playground-v2.5 [22]. In this scenario, classifiers are tasked not only with identifying real versus generated distribution but also with determining which specific generated distribution an image comes from. Notably, we find that in all settings, classifiers achieve $>76\%$ distribution classification accuracy. Meanwhile, we also observe significant improvements as the number of training samples increases from 5k to 10k, which aligns with the label-to-image scenario.
5.2 Self-supervised classifiers can identify the generated images
Except for the supervised learning protocol, we also explore the self-supervised learning protocol. Specifically, we use the self-supervised pre-trained models, such as MAE [55] and MoCo v3 [56], to extract image features and then train a linear classifier on these features. Following, we use the self-supervised classifier to refer to this linear classifier. We report the distribution classification accuracy of self-supervised classifiers in Tab. 3. We observe two intriguing phenomena: First, distinct from semantic classification results on ImageNet, in the distribution classification task, masked image modelling-based MAE is more effective at identifying generated images related to ImageNet compared to the contrastive learning-based MoCo v3. Second, the self-supervised classifiers achieve relatively high accuracy in tasks related to ImageNet. However, there is a notable decline in accuracy when these models are transferred to CIFAR-10. We hypothesize that this occurs for two main reasons: Firstly, MAE and MoCo v3 are pre-trained specifically on ImageNet, not on CIFAR-10. Secondly, diffusion models exhibit high performance on CIFAR-10 image generation, and the mere use of a linear classifier is insufficient for distinguishing between real and generated images.
Table 3: Linear classifier on self-supervised features achieves high accuracy on ImageNet-related combinations, where the model is pre-trained, but accuracy declines when transferred to CIFAR-10. We use ViT-B as the backbone.
Real dataset Generative model Self-supervised method Accuracy
C U [20] MAE 71.83
MoCo v3 73.39
C E [26] MAE 64.83
MoCo v3 67.69
I-256 U-H/2 [20] MAE 91.65
MoCo v3 77.38
I-512 E2-XXL [19] MAE 81.80
MoCo v3 74.30
Conclusion 1.
Despite the remarkable performance of diffusion models in generating realistic images, classifiers easily distinguish real from generated images with far fewer training samples than required for the diffusion models, achieving high accuracy across various combinations. Besides, self-supervised classifiers also demonstrate the ability to identify generated images.
6 Classifier contradictions with FID and human judgments
Previously, we demonstrated that classifiers can effortlessly distinguish between real and generated images. In this section, we explore the intriguing contradictions between classifiers and both FID and human judgments through two sets of experiments. The first experiment involves EDM2-XS and EDM2-XXL, which share identical network architectures and training methodologies but differ in the number of parameters, resulting in FID scores of 2.91 and 1.81, respectively. The second experiment includes DiT-XL, U-ViT-H/2, and EDM2-XXL, which, despite having similar FID scores, differ in network architecture, sampling methods, and training protocols.
Table 4: Distribution classification accuracy on combinations of different diffusion models with similar FIDs. An intriguing phenomenon occurs when more distributions are added: the classifier achieves higher distribution classification accuracy.
Combinations Model Accuracy
U-H/2, D ResNet-50 99.66
ConvNeXt-T 97.59
ViT-S 86.13
U-H/2, D, I-256 ResNet-50 99.87
ConvNeXt-T 99.77
ViT-S 95.90
U-H/2, D, I-256, E2-XXL ResNet-50 99.91
ConvNeXt-T 99.92
ViT-S 98.13
Figure 3: User study. Classifiers can distinguish between different diffusion models with similar FIDs but struggle with variations of the same diffusion model differing in parameters. In contrast, humans show the opposite tendency.
6.1 User study
Here, we conduct a user study to explore human capabilities in distinguishing generated distributions from real ones and classifying from which diffusion model an image is generated (see Fig. 10). Details of the study can be found in Appendix C.4. The results are presented in Fig. 3, with classifier performance represented by the median accuracy achieved using three architectures. Specifically, humans are unable to distinguish between different diffusion models with similar FIDs, yet they demonstrate some capability in distinguishing different variants of EDM2.
6.2 Classifiers distinguish between different diffusion models with similar FIDs
As illustrated in Tab. 4, we present the distribution classification accuracy for three combinations of datasets generated by diffusion models with similar FIDs. We observe that classifiers can effortlessly distinguish between different diffusion models, even those with closely matched FID scores, while humans cannot. Notably, an intriguing phenomenon occurs when more datasets are added: the classifier achieves even higher classification accuracy. This phenomenon differs from the findings in  [28]. A possible explanation is that all these datasets are related to ImageNet, whether directly from ImageNet or from diffusion models trained on it. Thus, adding new distributions enables the classifier to apply newly learned patterns to enhance the classification accuracy of previous combinations.
6.3 Classifiers struggle to distinguish different models in the same family
As shown in Tab. 5, we present the distribution classification accuracy for combinations of EDM2-XS and EDM2-XXL, which are the smallest and largest models in the EDM2 family, respectively. We observe that classifiers struggle to distinguish between different models within the same family. Only the ResNet-50 demonstrates some ability to classify them, while the other two models perform close to random selection. Notably, humans also show some ability to classify them (see Fig. 3).
Conclusion 2.
Classifiers conflict with FID and human assessments in distribution classification. They can distinguish between different diffusion models with similar FID but different inductive biases but struggle to do so within the same family, where models share inductive biases but differ in visual quality. In contrast, humans show a different trend.
7 How classifiers distinguish between different distributions
The previous findings illustrate the classifier’s capability to differentiate between real and generated images and detect different inductive biases in diffusion models. In this section, we delve into how classifiers distinguish between different distributions to explore the reasons behind this capability.
7.1 Classifier uses edge information to distinguish real from generated images
We first use Grad-CAM [31] to investigate which aspects of image classifiers depend on to distinguish between real and generated images. We use the combination of U-ViT-H/2 and ImageNet-256, with ResNet-50 as the classifier architecture. We observe that classifiers rely mainly on edge information of an image to make their classification, regardless of whether the image is real or generated (see Fig. 4, first column). This led us to question whether the classifier’s ability to distinguish between real and generated images would persist if edge information were reduced. To validate this, we center-crop images from a resolution of 256 to smaller sizes of 128, 64, 32, and 16, respectively, and train the classifier from scratch. The results, shown in Tab. 6, indicate that even without original edge information, the classifier can also distinguish between real and generated images with high accuracy based on the new edge information presented at each lower resolution (see Fig. 4).
Table 5: Classifiers struggle to differentiate between the smallest and largest models of EDM2. Despite notable differences in FID, the classifiers show limited ability to accurately distinguish between different EDM2 variants.
Combinations Classifier Accuracy (%)
E2-XS, E2-XXL ResNet-50 74.91
ConvNeXt-T 57.35
ViT-S 59.57
Table 6: Classifiers achieve high accuracy across crop sizes. denotes center crop. denotes random crop, where results are averaged over four trials.
Combinations Crop size Accuracy (%)
U-H/2, I-256 None 99.96
128 99.96
64 99.97
32 99.97
16 99.78
128 99.97
64 99.97
Figure 4: Grad-CAM [31] results on validation images (top) and samples of U-ViT-H/2 (bottom).
7.2 Classifiers can discern real from generated images using any part
Building on our findings that classifiers can effectively distinguish between real and generated images using just the central portion, we extended our investigation to see if classifiers can make distinctions using any arbitrary part of an image. To explore this, we focused on two resolutions, 128 and 64, and for each resolution, we randomly cropped images four times to train the classifiers (see Fig. 7 in the Appendix for visualization of crops). For images cropped to the resolution of 128 and 64, classifiers achieve an accuracy of $99.97\pm 0.004$ and $99.97\pm 0.007$, respectively (see Tab. 6). The quantitative results on Tab. 6 demonstrate that, to a certain extent, classifiers are capable of classifying whether an image is real or generated based on any part of the image.
(a) Accuracy vs. low-frequency filter threshold
(b) Accuracy vs. high-frequency filter threshold
(c) Accuracy vs. band-frequency filter threshold
Figure 5: Classifier achieves high accuracy across various combinations with limited frequency components, except when only a little low-frequency components is present.
7.3 The impact of frequency on distribution classification task
Having established that classifiers mainly rely on edge information to classify both original and cropped images, regardless of the position and size of the cropped images, we will next explore the impact of frequency on classification. To this end, we preprocess the images with low-pass, high-pass and band-pass filters as defined in Sec. 2.2. For low-pass and high-pass filters, we apply four thresholds: 10, 30, 50 and 100. For the band-pass filter, we choose four threshold intervals: 0-10, 10-30, 30-50, and 50-100. We visualize an example of original and processed images in Fig. 6.
Figure 6: Visualization of frequency domain processing. Top: Low-Pass, Middle: High-Pass, each processed with increasing thresholds: 10, 30, 50, 100. Bottom: Band-Pass, processed with band thresholds: 0-10, 10-30, 30-50, 50-100.
As shown in Fig. 5, using ResNet-50 as the classifier architecture, we observe two intriguing phenomena when classifying images as either generated or real, based on different frequency filters. First, classifiers can effortlessly classify images even when primarily high-frequency components are present (i.e., threshold at 100). Second, classifiers can classify images even with any interval of frequency components, except when only very little low-frequency information is available.
Moreover, we also conduct an experiment on the combinations of EDM2-XS and EDM2-XXL (see details in Appendix. C.3), we observe that for models within the same diffusion model family, which share inductive biases but differ in visual quality, classifiers unable to distinguish between them based on any specific frequency band, achieving accuracy close to random guessing (around 50%).
Conclusion 3.
Classifiers primarily rely on edge information to distinguish between real and generated images. Even when original edges are removed via center cropping, they utilize new edges at lower resolutions. Additionally, classifiers can distinguish based on any part and maintain high accuracy unless only minimal low-frequency information is available. In contrast, for models within the same diffusion family, classifiers’ accuracy drops close to random guessing (around 50%) when distinguishing based on any specific frequency bands.
8 Conclusions
Our paper proposes the distribution classification tasks to explore a fundamental question: How far is the generated distribution from the real distribution? Our study reveals that classifiers can effectively distinguish between diffusion-generated and real distributions, despite the low FID scores and lifelike appearance of the generated images. Moreover, classifiers are more sensitive to the inductive biases of different diffusion models than humans, who rely more on visual quality. Finally, classifiers primarily use edge information for these distinctions, maintaining high accuracy even with limited frequency components, except when only minimal low-frequency components are present.
Limitations and social impact. Our findings are based on specific datasets and model architectures, which may limit their generalizability. Future work can explore a wider range of datasets to validate our results. Further, our paper may draw attention to enhancing the edge or high-frequency details in generated images in the future, potentially making "DeepFakes" more difficult to detect.
Appendix A Detail experiment settings
We implement our experiments upon the official code of ConvNeXt [53], U-ViT [20], DiT [21], EDM [26], EDM2 [19]. We also utilize the official models of Pixart-$\alpha$ [23], SDXL [24], and Playground-v2.5 [22]. The respective links and licenses are detailed in Tab. 7.
Table 7: Code links and licenses.
Method Link License
ConvNeXt https://github.com/facebookresearch/ConvNeXt MIT License
U-ViT https://github.com/baofff/U-ViT MIT License
DiT https://github.com/facebookresearch/DiT CC BY-NC 4.0
EDM https://github.com/NVlabs/edm CC BY-NC-SA 4.0
EDM2 https://github.com/NVlabs/edm2 CC BY-NC-SA 4.0
Model Link (add ‘https://huggingface.co/’) License
PixArt-$\alpha$ PixArt-alpha/PixArt-XL-2-1024-MS Open RAIL++-M
SDXL stabilityai/stable-diffusion-xl-base-1.0 Open RAIL++-M
Playground-v2.5 playgroundai/playground-v2.5-1024px-aesthetic Playground v2.5
We present the main experiment settings as follows.
Dataset. For label-to-image, we consider the CIFAR [48] and ImageNet [49] datasets, which are well-established and widely recognized benchmarks in the field of image generation. For text-to-image, we utilize the COCO2014 dataset [50], known for its rich annotations and diverse image content. For CIFAR-10, we use the real CIFAR-10 training set and validation set to construct the real distribution and use the diffusion model to generate 50k training images and 10k validation images to construct the generated distribution. In the case of ImageNet, we consider 256 and 512 resolutions, which are common resolutions for ImageNet image generation tasks. We randomly sample 100k training images and 50k validation images from the training set and the validation set of ImageNet to construct the real distribution and use the diffusion models to generate 100k training images and 50k validation images to construct the generated distribution. For the real dataset, we adopt the data processing method from ADM [51] to modify the ImageNet dataset into two common resolutions: ImageNet-256 and ImageNet-512, where the numbers indicate the resolution of the data. In addition, for the COCO2014 dataset, by default, we construct the real distribution by randomly sampling 10k training images and 1k validation images from the respective training and validation sets of COCO2014. Each image in this dataset is associated with five captions. To create the generated distribution, we randomly select one of the five captions for each real image to serve as a prompt. These prompts are then used by the diffusion models to generate an equivalent number of images to construct the generated distribution.
Classifier. By default, we employ the ResNet-50 [52] as the classifier architecture. For completeness, we also consider ConvNeXt-T [53] and ViT-S [54]. Our pre-processing protocol follows the standard supervised training approach [53]. Specifically, during training, classifiers process randomly augmented crops of $224\times 224$ images. During validation, images are resized so that their smaller dimension reaches 256 pixels while preserving the original aspect ratio. Subsequently, these images are center cropped to $224\times 224$ pixels before being fed into the model. For the experiments on CIFAR-10, we initialize our classifier with the ResNet-50, pre-trained on ImageNet. For completeness, we also present the result trained from scratch (see experiments in Appendix. C.1). In the case of ImageNet, we opt to train the ResNet-50 model from scratch.
Diffusion model. For the generation of CIFAR-10, we consider two diffusion models: EDM [26] and U-ViT [20]. Both models have demonstrated strong generation performance on the CIFAR-10 [48]. Quantitatively, EDM achieves an FID of 1.79, and U-ViT achieves an FID of 3.11. To improve efficiency, we modified U-ViT’s sampling method from Euler-Maruyama to DPM-Solver [57] and reduced the sampling steps from 1,000 to 50. These adjustments resulted in U-ViT achieving an FID of 3.65 on CIFAR-10. For the generation of ImageNet-256, we explore three diffusion models: EDM2 [19], U-ViT-H/2 [20], DiT-XL [21]. Both DiT and U-ViT are prominent diffusion transformer architectures known for their scalability and strong performance. U-ViT-H/2 achieves an FID of 2.29 on ImageNet-256, and DiT-XL/2 achieves an FID of 2.27. We consider EDM2 to incorporate a UNet-based architecture, which was traditionally used before the rise of diffusion transformers. Since EDM2 is originally designed for ImageNet-512 generation, we resize the generated images from 512 to 256 resolution to suit our ImageNet-256 task. In this way, EDM2-XXL achieves an FID of 2.14 on this task, which is similar to the FID achieved by U-ViT and DiT. For the generation of ImageNet-512, we use EDM2 [19], which achieves state-of-the-art performance on this task with an FID of 1.81. For the generation of COCO, we consider three state-of-the-art text-to-image diffusion models: Pixart-$\alpha$ [23], SDXL [24], and Playground-v2.5 [22].
Evaluation. We use the top-1 accuracy on the validation set to evaluate classification performance.
Training settings. The complete training settings of ResNet-50 are reported in Tab. 8 for combinations related to CIFAR-10 and Tab. 9 for combinations related to ImageNet.
Table 8: Training settings for CIFAR-10.
Config Value
Optimizer AdamW
Learning rate 4e-4
Weight decay 0.05
Optimizer momentum $\beta_{1},\beta_{2}$=0.9, 0.999
Batch size 256
Learning rate schedule Cosine decay
Warmup epochs 0
Training epochs 50
Augmentation RandAug (9, 0.5) [58]
Label smoothing 0.1
Mixup [59] 0
Cutmix [60] 0
Table 9: Training settings for ImageNet.
Config Value
Optimizer AdamW
Learning rate 1e-3
Weight decay 0.3
Optimizer momentum $\beta_{1},\beta_{2}$=0.9, 0.95
Batch size 4096
Learning rate schedule Cosine decay
Warmup epochs 20
Training epochs 200
Augmentation RandAug (9, 0.5) [58]
Label smoothing 0.1
Mixup [59] 0.8
Cutmix [60] 1.0
Appendix B Computational cost
Our experiments were conducted on RTX 3090 and V100 GPUs. The detailed computational costs are presented in Tab. 10. Training epochs were set to 50 for CIFAR-10 and 200 for ImageNet. The number of epochs trained on CIFAR-10 is relatively low because we use a pre-trained model to initialize our classifier, enabling faster convergence.
Table 10: Training time of classifiers.
Model Combinations Epochs GPU-type GPU-nums Hours
ResNet-50 C, U 50 3090 4 2
ViT-S C, U 50 3090 4 2
ConvNeXt-T C, U 50 3090 4 2
ResNet-50 I-256, U-H/2 200 V100 8 6
ViT-S I-256, U-H/2 200 V100 8 9
ConvNeXt-T I-256, U-H/2 200 V100 8 8
ResNet-50 I-512, E2-XXL 200 V100 8 9
ViT-S I-512, E2-XXL 200 V100 8 11
ConvNeXt-T I-512, E2-XXL 200 V100 8 10
Appendix C Additional results
C.1 Results of CIFAR-10
In order to ensure the completeness of the experiment, we are here to present the result of CIFAR-10 trained from scratch. We present the results in Tab. 11. If there is no prior knowledge, classifiers struggle to distinguish between real and generated data in datasets with low resolution, such as CIFAR-10 (i.e., 32x32). However, the use of a pre-trained model allows the features learned from the ImageNet dataset to aid in differentiating between real and generated images in CIFAR-10.
Table 11: Distribution classification accuracy on CIFAR-10. "Pretrained" indicates that the classifier was initialized with a model pretrained on ImageNet, while "Scratch" indicates that the classifier was trained from scratch.
Real dataset Generative model Model Pretrained Scratch
C E [26] ResNet-50 96.25 56.13
ViT-S 89.38 55.32
ConvNeXt-T 98.43 53.27
C U [20] ResNet-50 99.92 56.70
ViT-S 98.04 52.66
ConvNeXt-T 99.96 56.37
C.2 Visualization of crops
As shown in Fig. 7, we present the visualization of crops mentioned in Sec. 7.1 and Sec. 7.2.
Figure 7: Visualization of crops. With each resolution represented by a different color.
C.3 Frequency analysis on the combination of EDM2-XS and EDM2-XXL
We preprocess the generated images using band-pass filters as defined in Sec. 2.2, with four threshold intervals: 0-10, 10-30, 30-50, and 50-100. An example of original and processed EDM2-XXL generated images is shown in Fig. 8. As shown in Tab. 12 and Fig. 9, for EDM2-XS and EDM2-XXL, the smallest and largest models in the EDM2 family, classifiers’ accuracy approaches random guessing (around 50%) across different threshold intervals. This indicates that for models within the same diffusion model family, which share inductive biases but differ in visual quality, classifiers unable to distinguish between them based on any specific frequency band.
Figure 8: Visualization of frequency domain processing of EDM2-XXL. The image shows the results after applying a band-pass filter to EDM2-XXL with band thresholds of 0-10, 10-30, 30-50, and 50-100, from left to right.
Table 12: Classification accuracy of ResNet-50 on EDM2-XS and EDM2-XXL combinations after applying band-pass filters.
Combinations Classifier Threshold Intervals Accuracy
E2-XS, E2-XXL ResNet-50 0-10 57.94
10-30 58.96
30-50 58.12
50-100 50.48
Figure 9: Accuracy vs. band-frequency filter threshold
C.4 User study
Fig. 10 shows a screenshot of the interaction interface used in our user study. There are eleven participants. We design three groups of experiments, in each of which the participants are asked to classify 32 pairs of images. The first set involved distinguishing generated images from real ones, where the real images are from ImageNet-256, and the generated images are from U-ViT-H/2. The second set involved classifying two diffusion models with similar performance, specifically DiT-XL/2 and U-ViT-H/2. The final set included images from EDM2-XS and EDM2-XXL, which utilize the same training methodology but differ in the number of parameters, leading to different FID. For the first set, participants were required to identify the real images. In the second set, they were tasked with identifying DiT images; since they lacked prior exposure to these images, we provided some reference images from DiT and U-ViT during the test. In the third experiment, participants were asked to determine which images were of better quality.
(a) I-256 vs. U-ViT-H/2
(b) U-ViT-H/2 vs. D
(c) E2-XS vs. E2-XXL
Figure 10: Screenshot of user study. Participants are asked to distinguish generated distributions from real ones and to classify which diffusion model generated a given image. Each group of experiments is illustrated with a separate example here. In each set of experiments, we also randomized the order to prevent examples from the same set from influencing each other.